In 8: 00 AM, the population of a bacteria is 1000. At 11: 30 AM, the number of bacteria triples. At what time will the population become 300 times the initial population of bacteria?

1
Expert's answer
2021-12-20T16:15:03-0500

Initial population of bacteria (at 8:00 AM) =1000

300 times the initial population of bacteria = 1000*300 = 300,000

Number of bacteria tripled at 11:30 AM = 3.5 hours

The exponential form is;

"P(t)=P_0e^{rt}"


"300000=1000*3^{(\frac{t}{210})}"


"ln(300)={(\frac{t}{210})}\ln3"


"{(\frac{t}{210})}={(\frac{ln300}{ln3})}"


"t={(\frac{ln300}{ln3})}\times210"


"t=1090.28\approx1090 mins"


t = 18 hours 10 mins


"Time=8:00 AM+18hours10mins=2:10AM"